OpenStudy (anonymous):

Solve the quadratic equation by completing the square. -3x^2+9x=1

OpenStudy (anonymous):

This is my answer.

OpenStudy (solomonzelman):

-3x^2+9x=1 3x^2-9x=-1 x^2-3x=-1/3 x^2-3x=-1/3 (x-1/2)^2- 5/2=-1/3 (x-1/2)^2=-1/3+5/2 (x-1/2)^2=13/6 x-1/2= +-square root (13/6) x= +-square root (13/6)+ x= +-square root (78) /6 + 1/2  I think that is the answer.